OpenStudy (magbak):

I will award a medal!!!!!!!!!!!!!!!!!!!!!!!!!!!!!!!!! Please help!!!!!!!!!!!!!!!!!!!!!!!!!!!!!!!!!!!!!!!!! Using complete sentences, explain how to find the zeros of the function f(x) = 2x3 – 9x + 3. Create your own polynomial with a degree greater than 2 and find the zeros of the function.

OpenStudy (magbak):

Thank you for showing up.

OpenStudy (anonymous):

Well first you could use an online graqphing tool

OpenStudy (whpalmer4):

Personally, I would type \[\text{Solve}\left[2x^3-9x+3==0,x\right]\]into my handy Mathematica window :-)

OpenStudy (magbak):

What is your handy mathematical window

OpenStudy (anonymous):

It's a computer program

OpenStudy (whpalmer4):

Mathematica is a math workbench if you will, built by the people who built WolframAlpha

OpenStudy (magbak):

OHHHHHHHHHHHHHH so I went on it and it did not give me the zeros.

OpenStudy (anonymous):

the website i gave you? its sohuld be right below the grpah of it

OpenStudy (whpalmer4):

it did, but you might not have understood it :-)

OpenStudy (magbak):

Probably.:)

OpenStudy (whpalmer4):

it calls them roots, not zeros...

OpenStudy (magbak):

Oh so they are not whole numbers ?

OpenStudy (whpalmer4):

for some amusement, click the "Exact forms" button

OpenStudy (anonymous):

nope, because you could factor it easily if so

OpenStudy (magbak):

Ok so what do I put in as a answer or do I just leave it blank.

OpenStudy (whpalmer4):

Oh, you could mumble something about using Cardano's solution for the depressed cubic, that always impresses the girls :-)

OpenStudy (anonymous):

You can use Descartes' rule of signs to help with the possible number of real zeros, but that will be of limited help. Much more importantly, you can try all positive and negative factors of 3 over all possible factors of 2, but that will only help find rational zeros. There are some much more difficult techniques that are rarely even taught anymore variously called either Cardano or Tartaglia techniques (among others) that will give you your zeros, but the techniques are extremely complicated and beyond the high-school level. Hardly even taught in college anymore.

OpenStudy (magbak):

OK so my teacher is a guy and I do not want a academic integrity case.

OpenStudy (anonymous):

This will give you a glimpse of the complexity of the techniques: http://en.wikipedia.org/wiki/Cubic_function

OpenStudy (whpalmer4):

oh, no, I mean the girls in the library who are watching you do the homework, not the teacher :-) start trying values of x from say 0, moving in each direction by 1. so 0, 1 2, 3 4 5 notice when the sign of f(x) changes. that means there's a zero between the current value of x and the previous one. also do that going in the negative direction along the axis.

OpenStudy (magbak):

Ok so I just use the dcartas rule and that is it right.

OpenStudy (magbak):

I do my home work at home and I am cool like that. Thnx

OpenStudy (whpalmer4):

no, Descartes just tells you possible combinations of types of roots/zeros. to actually find zeros, you need either heavy-duty math, or an iterative numeric solver, or use my approach of trying sample values...

OpenStudy (anonymous):

Descartes will only tell you if there is exactly one negative or exactly one positive real root. Say there is one positive real root. You could have zero or two negative real roots.

OpenStudy (magbak):

Ok so I get both of your explanations but after that what do I do for the second question.

OpenStudy (anonymous):

You can work backwards with: y = (x - a)(x - b)(x - c) Pick your a, b, and c values and make your polynomial.

OpenStudy (anonymous):

I loathe Descartes' rule and hate it even more when I get x^5-4x^3+8x^2-6x-2 and (3-i) is a root and they say to factor the whole thing -__-

OpenStudy (magbak):

Ok say I pick a= 2,b=4,c=6 now can I get help finding the zeros.

OpenStudy (anonymous):

If you select your a, b, and c, make your polynomial, you can demonstrate the solving by synthetic division. That's what I would do. ok, y= (x - 2)(x - 4)(x - 6) Now, multiply out.

OpenStudy (anonymous):

your roots are the numbers you pick they are the factor of the polynomial, where the graph touches zero

OpenStudy (anonymous):

y = (x^2 - 6x + 8)(x - 6) y = x^3 - 12x^2 + 44x - 48

OpenStudy (magbak):

Ok so routs are the location where the graph touches zero. Also @tcarroll010 I was getting to the equation and it resulted in y=x^3 + 32x -48

OpenStudy (anonymous):

Now, you can solve this by synthetic division and you will get your roots. You know it will work, because you "pre-selected" them.

OpenStudy (magbak):

But why divide if I already know it is right and also if I am only asked for the roots.

OpenStudy (anonymous):

Well, you would either use synthetic division or long division because your part b asks you to show how to solve it.

OpenStudy (anonymous):

Oh yeah true you just gotta show ok this is how I made it, and this i s how i break it down

OpenStudy (magbak):

No you missed up the first question and the second the second only asks for the equation and the roots the first one asks for the explanation.

OpenStudy (anonymous):

You see, the part about pre-picking your zeros is not part of what you show others. It is just to come up with something you know is solvable. You don't want to arbitrarily pick coefficients and come up with some impossibly-hard cubic that requires Tartaglia or Cardano.

OpenStudy (anonymous):

Your part b, where there is a sentence starting out with "create your own" asks you to solve it. So you want to pick good coefficients, which means you better work backwards like I outlined.

OpenStudy (magbak):

Ok that is one part and for the roots what do I do.

OpenStudy (anonymous):

As for your first question, my first several posts were aimed solely at that particular question.

OpenStudy (magbak):

no I mean the roots for the second question sorry if I am giving you a hard time.

OpenStudy (anonymous):

That is described in my post which starts out: "If you select . . ." that post and the one after it.

OpenStudy (magbak):

Ok so my roots are 2,4,6 am I correct or not?

OpenStudy (anonymous):

Yes, those are the roots I used to build you your cubic polynomial.

OpenStudy (magbak):

Why thank you so very much I really appreciate your help and the others and your kind patience. THANK YOU A+++++++++++++ HELP REALY FIND YOU TERRIFIC AT EXPLAINING AND YOU ARE A OCEAN OF KNOWLEDGE

OpenStudy (anonymous):

You are very appreciative! You're great to work with!